# Perturbation-Robust Predictive Modeling of Social Effects by Network Subspace Generalized Linear Models

### v4

@@ Appendix D Proof of Lemma 3
$\left\|T_{n}^{\top}\tilde{g}_{i}\right\|\leq\left\|T_{n}^{\top}\tilde{g}_{i}-g_{i}\right\|+\|g_{i}\|\leq C_{2}n^{1/2}\bar{\tau}_{n}+C\leq 2C,$ (36)
@@
$\big{\|}T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F({{\gamma}})\big{\|}\leq\Psi(\|\gamma\|)\bar{\tau}_{n},$
@@
$1/C-\Psi(\|\gamma\|)\bar{\tau}_{n}\leq\lambda_{\min}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq\lambda_{\max}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq C+\Psi(\|\gamma\|)\bar{\tau}_{n}.$ (37)
@@
$\leq n^{-1}\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\sum_{i=1}^{n}\left(C_{2}\bar{\tau}_{n}+2C\right)\left\|T_{n}^{\top}\tilde{g}_{i}-g_{i}\right\|$
@@
$\leq\left(\left(2C+C_{2}\right)\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\right)C_{2}\bar{\tau}_{n}.$
@@
$\leq C^{2}C_{2}\|\gamma\|\max_{1\leq i\leq n}|\varphi^{\prime}(p^{\top}_{i}\gamma)|\bar{\tau}_{n},$
@@
$\left\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\right\|$ $\leq\left(C^{3}\|\gamma\|\max_{1\leq i\leq n}|\varphi^{\prime}(p^{\top}_{i}\gamma)|+3C^{2}\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\right)\bar{\tau}_{n}$
$\leq\left(C^{2}C_{2}\|\gamma\|\max_{\left|t\right|\leq 2C\|\gamma\|}|\varphi^{\prime}\left(t\right)|+\left(2C+C_{2}\right)C_{2}\max_{\left|t\right|\leq 2C\|\gamma\|}|\varphi\left(t\right)|\right)\bar{\tau}_{n}.$
@@
$\lambda_{\min}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})$ $\geq\lambda_{\min}(F(\gamma))-\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\|\geq 1/C-\Psi(\|\gamma\|)\bar{\tau}_{n}.$
@@
$\lambda_{\max}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq\lambda_{\max}(F(\gamma))+\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\|\leq C+\Psi(\|\gamma\|)\bar{\tau}_{n}.$
@@
Since it is assumed that $\bar{\tau}_{n}\to 0$, this implies $\Psi(\left\|\gamma^{*}\right\|)\bar{\tau}_{n}$ is close to zero as $n$ is sufficiently large, and (37) implies (38). ∎
@@
$\leq C_{2}\bar{\tau}_{n}\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|h\left(t\right)|\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq 2C\left\|\gamma^{*}\right\|}|\eta^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\eta\left(t\right)|\right).$
@@
$\left\|B_{1}\right\|\leq C_{2}\left(6C^{2}+1\right)M^{2}\bar{\tau}_{n}=o(n^{-1/2}).$ (41)
@@
$\leq C_{2}\bar{\tau}_{n}\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq 2C\left\|\gamma^{*}\right\|}|\left(\eta\cdot h\right)^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\left(\eta\cdot h\right)\left(t\right)|\right)$
$\leq C_{2}\left(6C^{2}+1\right)M\bar{\tau}_{n}=o(n^{-1/2}).$
@@
$\sum_{i=1}^{n}|a_{i}|$ $\leq n^{1/2}C_{2}\bar{\tau}_{n}\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq n^{3/2}2C\left\|\gamma^{*}\right\|}|\eta^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\eta\left(t\right)|\right)$
$\leq n^{3/2}C_{2}\left(6C^{2}+1\right)M\bar{\tau}_{n}=o(n).$
@@
Note that $\Phi_{1}$ is the expectation of the second term on the right-hand side of (44), while $-\Phi_{2}-\Phi_{3}$ is its centered version, where $\Phi_{2}$ does not depend on $\gamma$ and $\Phi_{3}$ depends on $\gamma$. We will bound $\Phi_{1},\Phi_{2}$, and $\Phi_{3}$ separately. Regarding $\Phi_{1}$, by (46) and the triangle inequality,
@@
$\left\|T_{n}^{\top}\tilde{g}_{i}-{g}_{i}\right\|\leq C_{2}n^{1/2}\bar{\tau}_{n}=o(1),\quad\|T^{-1}_{n}\gamma\|\leq 2\left\|\gamma^{*}\right\|.$
@@
$\left\|\bar{g}_{i}\right\|=\left\|{g}_{i}+\bar{t}_{i}(T^{\top}_{n}\tilde{g}_{i}-g_{i})\right\|\leq\left\|g_{i}\right\|+\bar{t}_{i}\left\|T^{\top}_{n}\tilde{g}_{i}-g_{i}\right\|\leq C+\bar{t}_{i}C_{2}n^{1/2}\bar{\tau}_{n}\leq 2C.$
@@
[…]
@@
[…]
@@
[…]
@@
It remains to bound $\|S(\gamma^{*})\|$. Note that it follows directly from the definition of $S(\gamma^{*})$ in (12) that $\mathbb{E}[S\left(\gamma^{*}\right)]=0$ and $\operatorname{Cov}(S\left(\gamma^{*}\right))=n^{-1}F\left(\gamma^{*}\right)$ is a square matrix of size $(p+K-r)$. Therefore, by Markov inequality and Assumption 4, for any $t>0$,
@@ Appendix E The Proof of Theorem 3
[…]
@@
In light of Lemma 8 and particularly the bound (D) in its proof, we will approximate $\mathbb{H}(\hat{\gamma})$ by $T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n}$, and in turn by $F(\gamma^{*})T_{n}$. Accordingly, we decompose the expression above as
@@
These imply, in particular, that $\|\mathbb{H}^{-1}(\hat{\gamma})\|$ and $\|(T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n})^{-1}\|$ are both bounded by some constant due to the continuity of the inverse map away from zero. Therefore,
@@
$\left\|\left(T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n}\right)-F(\gamma^{*})\right\|\leq\Psi(\|\gamma^{*}\|)\bar{\tau}_{n}=o(n^{-1/2}).$ (49)
@@
For each positive integer $n$, let $X_{nj}$, $j=1,2,...,n$, be independent random variables with $\mathbb{E}[X_{nj}]=0$ and $\mathbb{E}[X_{nj}^{2}]=\sigma_{nj}^{2}<\infty$. Denote $B_{n}^{2}=\sum_{j=1}^{n}\sigma_{nj}^{2}$ and assume that for each $\varepsilon>0$,
@@
$z_{1:r}=\frac{n^{-1}{Z}_{1:r}^{T}\hat{\mathcal{P}}_{R}XGu}{\|n^{-1}{Z}_{1:r}^{T}\hat{\mathcal{P}}_{R}XGu\|},\qquad z_{(r+1):(K+p-r)}=0.$
@@
$n^{-1}\left\|{Z}_{1:r}^{T}\hat{\mathcal{P}}_{R}XGu\right\|$$\geq$ $n^{-1}\left\|{Z}_{1:r}^{T}XGu\right\|-n^{-1}\left\|{Z}_{1:r}^{T}(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})XGu\right\|$
$\geq$ $c-n^{-1}\left\|{Z}_{1:r}^{T}\right\|\left\|\hat{\mathcal{P}}_{R}-\mathcal{P}_{R}\right\|\left\|XGu\right\|$
@@
$z_{1:r}=0,\qquad z_{(r+1):p}=\frac{n^{-1}{Z}_{(r+1):p}^{T}\hat{\mathcal{P}}_{R}XGu}{\|n^{-1}{Z}_{(r+1):p}^{T}\hat{\mathcal{P}}_{R}XGu\|},\qquad z_{(p+1):(K+p-r)}=0.$
@@
$n^{-1}\left\|{Z}_{(r+1):p}^{T}\hat{\mathcal{P}}_{C}XGu\right\|\geq c-C^{1/2}C_{2}\bar{\tau}_{n}>c/2.$
@@
By the Schur complement formula, $\tilde{F}^{-1}_{3}(\hat{\gamma})$ equals
@@
$\kappa(\hat{\gamma})=\operatorname{diag}\left((h^{\prime}(\tilde{g}^{\top}_{i}\hat{\gamma}))^{2}/v(\tilde{g}^{\top}_{i}\hat{\gamma})\right).$
@@
$n\left(\hat{\alpha}-n^{-1}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)^{\mathrm{T}}\tilde{O}\left(\hat{\alpha}-n^{-1}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)\ \to\ \chi_{K-r}^{2},$
where $\tilde{O}=n^{-1}(\kappa(\hat{\gamma})-\kappa(\hat{\gamma})\tilde{Z}(\tilde{Z}^{\top}\kappa(\hat{\gamma})\tilde{Z})^{-1}\tilde{Z}^{\top}\kappa(\hat{\gamma}))$. The proof is complete. ∎
@@ Appendix F The Proof of Corollary 2
$\leq$ $n^{-1/2}\left\|{\mathcal{P}}_{N}Z_{1:r}x_{1:r}\right\|+n^{-1/2}\left\|{\mathcal{P}}_{N}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\bar{\tau}_{n}$
$=$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{1:r}x_{1:r}\right\|+n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\bar{\tau}_{n}$
$=$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\bar{\tau}_{n}$
$\leq$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}\right\|\left\|x_{(r+1):K}\right\|+C_{1}\bar{\tau}_{n}$
$\leq$ $\sigma_{r+1}\|x\|+C_{1}\bar{\tau}_{n}<1,$

### v6

@@ Appendix D Proof of Lemma 3
$\left\|T_{n}^{\top}\tilde{g}_{i}\right\|\leq\left\|T_{n}^{\top}\tilde{g}_{i}-g_{i}\right\|+\|g_{i}\|\leq C_{2}n^{1/2}\tau_{n}+C\leq 2C,$ (36)
@@
$\big{\|}T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F({{\gamma}})\big{\|}\leq\Psi(\|\gamma\|)\tau_{n},$
@@
$1/C-\Psi(\|\gamma\|)\tau_{n}\leq\lambda_{\min}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq\lambda_{\max}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq C+\Psi(\|\gamma\|)\tau_{n}.$ (37)
@@
$\leq n^{-1}\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\sum_{i=1}^{n}\left(C_{2}\tau_{n}+2C\right)\left\|T_{n}^{\top}\tilde{g}_{i}-g_{i}\right\|$
@@
$\leq\left(\left(2C+C_{2}\right)\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\right)C_{2}\tau_{n}.$
@@
$\leq C^{2}C_{2}\|\gamma\|\max_{1\leq i\leq n}|\varphi^{\prime}(p^{\top}_{i}\gamma)|\tau_{n},$
@@
$\left\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\right\|$ $\leq\left(C^{3}\|\gamma\|\max_{1\leq i\leq n}|\varphi^{\prime}(p^{\top}_{i}\gamma)|+3C^{2}\max_{1\leq i\leq n}\left|\varphi(\tilde{g}^{\top}_{i}T_{n}\gamma)\right|\right)\tau_{n}$
$\leq\left(C^{2}C_{2}\|\gamma\|\max_{\left|t\right|\leq 2C\|\gamma\|}|\varphi^{\prime}\left(t\right)|+\left(2C+C_{2}\right)C_{2}\max_{\left|t\right|\leq 2C\|\gamma\|}|\varphi\left(t\right)|\right)\tau_{n}.$
@@
$\lambda_{\min}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})$ $\geq\lambda_{\min}(F(\gamma))-\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\|\geq 1/C-\Psi(\|\gamma\|)\tau_{n}.$
@@
$\lambda_{\max}(T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n})\leq\lambda_{\max}(F(\gamma))+\|T_{n}^{\top}\tilde{F}(T_{n}\gamma)T_{n}-F(\gamma)\|\leq C+\Psi(\|\gamma\|)\tau_{n}.$
@@
Since it is assumed that $\tau_{n}\to 0$, this implies $\Psi(\left\|\gamma^{*}\right\|)\tau_{n}$ is close to zero as $n$ is sufficiently large, and (37) implies (38). ∎
@@
$\leq C_{2}\tau_{n}\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|h\left(t\right)|\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq 2C\left\|\gamma^{*}\right\|}|\eta^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\eta\left(t\right)|\right).$
@@
$\left\|B_{1}\right\|\leq C_{2}\left(6C^{2}+1\right)M^{2}\tau_{n}=o(n^{-1/2}).$ (41)
@@
$\leq C_{2}\tau_{n}\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq 2C\left\|\gamma^{*}\right\|}|\left(\eta\cdot h\right)^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\left(\eta\cdot h\right)\left(t\right)|\right)$
$\leq C_{2}\left(6C^{2}+1\right)M\tau_{n}=o(n^{-1/2}).$
@@
$\sum_{i=1}^{n}|a_{i}|$ $\leq n^{1/2}C_{2}\tau_{n}\left(2C\left\|\gamma^{*}\right\|\sup_{\left|t\right|\leq n^{3/2}2C\left\|\gamma^{*}\right\|}|\eta^{\prime}\left(t\right)|+\sup_{\left|t\right|\leq C\left\|\gamma^{*}\right\|}|\eta\left(t\right)|\right)$
$\leq n^{3/2}C_{2}\left(6C^{2}+1\right)M\tau_{n}=o(n).$
@@
Note that $\Phi_{1}$ is the expectation of the second term on the right-hand side of (44), while $\Phi_{2}-\Phi_{3}$ is its centered version, where $\Phi_{2}$ does not depend on $\gamma$ and $\Phi_{3}$ depends on $\gamma$. We will bound $\Phi_{1},\Phi_{2}$, and $\Phi_{3}$ separately. Regarding $\Phi_{1}$, by (46) and the triangle inequality,
@@
$\left\|T_{n}^{\top}\tilde{g}_{i}-{g}_{i}\right\|\leq C_{2}n^{1/2}\tau_{n}=o(1),\quad\|T^{-1}_{n}\gamma\|\leq 2\left\|\gamma^{*}\right\|.$
@@
$\left\|\bar{g}_{i}\right\|=\left\|{g}_{i}+\bar{t}_{i}(T^{\top}_{n}\tilde{g}_{i}-g_{i})\right\|\leq\left\|g_{i}\right\|+\bar{t}_{i}\left\|T^{\top}_{n}\tilde{g}_{i}-g_{i}\right\|\leq C+\bar{t}_{i}C_{2}n^{1/2}\tau_{n}\leq 2C.$
@@
$f(\beta)-f(\alpha)=\left(\int_{0}^{1}\frac{\partial f(\alpha+t(\beta-\alpha))}{\partial x}dt\right)(\beta-\alpha),$
@@
$\mathbb{H}(\gamma)=\int_{0}^{1}H\left(\gamma^{*}+t\left(T^{-1}_{n}\gamma-\gamma^{*}\right)\right)dt,\qquad H(\gamma)=T^{\top}_{n}(\partial\tilde{S}(\gamma)/\partial{\gamma}^{\top})T_{n}.$
@@
$\left\|\Phi_{2}\right\|$ $=$ $\left\|\int_{0}^{1}\left[H\left(\gamma^{*}+t\left(T^{-1}_{n}\gamma-\gamma^{*}\right)\right)-T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n}\right]dt\right\|$
$\leq$ $\int_{0}^{1}\left\|H\left(\gamma^{*}+t\left(T^{-1}_{n}\gamma-\gamma^{*}\right)\right)-T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n}\right\|dt$
@@
It remains to bound $\|S(\gamma^{*})\|$. Note that it follows directly from the definition of $S(\gamma^{*})$ in (12) that $\mathbb{E}[S\left(\gamma^{*}\right)]=0$ and $\operatorname{Cov}(S\left(\gamma^{*}\right))=n^{-1}F\left(\gamma^{*}\right)$ is a square matrix of size (p+Kr)(p+K-r). Therefore, by Markov inequality and Assumption 4, for any $t>0$,
@@ Appendix E The Proof of Theorem 3
$\mathbb{H}(\gamma)=\int_{0}^{1}H\left(\gamma^{*}+t\left(T^{-1}_{n}\gamma-\gamma^{*}\right)\right)dt,\qquad H(\gamma)=T^{\top}_{n}(\partial\tilde{S}(\gamma)/\partial{\gamma}^{\top})T_{n}.$
@@
[…]
@@
[…]
@@
$\left\|\left(T_{n}^{\top}\tilde{F}\left(T_{n}\gamma^{*}\right)T_{n}\right)-F(\gamma^{*})\right\|\leq\Psi(\|\gamma^{*}\|)\tau_{n}=o(n^{-1/2}).$ (49)
@@
For each positive integer $n$, let $X_{nj}$, j=1,2,,nj=1,2,...,n, be independent random variables with $\mathbb{E}[X_{nj}]=0$ and $\mathbb{E}[X_{nj}^{2}]=\sigma_{nj}^{2}<\infty$. Denote $B_{n}^{2}=\sum_{j=1}^{n}\sigma_{nj}^{2}$ and assume that for each $\varepsilon>0$,
@@
$z_{1:r}=\frac{n^{-1}{Z}_{1:r}^{\top}\hat{\mathcal{P}}_{R}XGu}{\|n^{-1}{Z}_{1:r}^{\top}\hat{\mathcal{P}}_{R}XGu\|},\qquad z_{(r+1):(K+p-r)}=0.$
@@
$n^{-1}\left\|{Z}_{1:r}^{\top}\hat{\mathcal{P}}_{R}XGu\right\|$ $\geq$ $n^{-1}\left\|{Z}_{1:r}^{\top}XGu\right\|-n^{-1}\left\|{Z}_{1:r}^{\top}(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})XGu\right\|$
$\geq$ $c-n^{-1}\left\|{Z}_{1:r}^{\top}\right\|\left\|\hat{\mathcal{P}}_{R}-\mathcal{P}_{R}\right\|\left\|XGu\right\|$
@@
$z_{1:r}=0,\qquad z_{(r+1):p}=\frac{n^{-1}{Z}_{(r+1):p}^{\top}\hat{\mathcal{P}}_{R}XGu}{\|n^{-1}{Z}_{(r+1):p}^{\top}\hat{\mathcal{P}}_{R}XGu\|},\qquad z_{(p+1):(K+p-r)}=0.$
@@
$n^{-1}\left\|{Z}_{(r+1):p}^{\top}\hat{\mathcal{P}}_{C}XGu\right\|\geq c-C^{1/2}C_{2}\tau_{n}>c/2.$
@@
[…]
@@
[…]
@@
$n\left(\hat{\alpha}-n^{-1}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)^{\top}\tilde{O}\left(\hat{\alpha}-n^{-1}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)\ \to\ \chi_{K-r}^{2},$
where $)\tilde{O}=n^{-1}(\kappa(\hat{\gamma})-\kappa(\hat{\gamma})\tilde{Z}(\tilde{Z}^{\top}\kappa(\hat{\gamma})\tilde{Z})^{-1}\tilde{Z}^{\top}\kappa(\hat{\gamma}$)). The proof is complete. ∎
@@ Appendix F The Proof of Corollary 2
$\leq$ $n^{-1/2}\left\|{\mathcal{P}}_{N}Z_{1:r}x_{1:r}\right\|+n^{-1/2}\left\|{\mathcal{P}}_{N}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\tau_{n}$
$=$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{1:r}x_{1:r}\right\|+n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\tau_{n}$
$=$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}x_{(r+1):K}\right\|+C_{1}\tau_{n}$
$\leq$ $n^{-1}\left\|W^{\top}_{(r+1):K}Z_{(r+1):K}\right\|\left\|x_{(r+1):K}\right\|+C_{1}\tau_{n}$
$\leq$ $\sigma_{r+1}\|x\|+C_{1}\tau_{n}<1,$